|
ID |
Date |
Author |
Subject |
|
80
|
Fri Sep 24 09:44:58 2021 |
JM ML | Vacuum gauge readout |
The analog readout of the Combivac controller for the IE514 ion gauge has been set up, readings are plotted to http://web-docs.gsi.de/~lestinsk/vacplot.php. Readings and plot should be updated every minute. |
|
79
|
Fri Sep 24 08:47:36 2021 |
JM RSS CB(Remotely) | Baking |
Baking halted on 23/9/21 16:16 as connection to TC08 module was lost. USB cable connection to TC08 module appears to have been damaged in shipment. USB hub and cable replaced, connection re-established and baking restarted 17:27.
Ramping of baking limited to 40C by the maximum difference between parts of the chamber being too high, this safety feature prevents high thermal gradients. Limitation caused by uneven heating due to placement of heating elements resulting in the gate valve end of the chamber being much cooler than the rest of chamber. The pressure increased to ~8e-7mbar at 40C due to increased outgassing. Baking stopped at 07:00 24/9/21, heating elements were repositioned and the baking jacket reassembled. Code recompiled to include the rate and temperatures limits from elog 76.
Baking restarted at 10:00 |
|
78
|
Fri Sep 24 07:04:57 2021 |
RSS JM | IE 514 pin arrangement and degas instructions |
Fabre (German) Color (English)
grau grey
blau blue
rot red
braun brown
weiss white
rosa pink
gelb/gruen yellow/green |
|
77
|
Thu Sep 23 14:19:27 2021 |
CB (remotely) | Checking baking remotely - instructions |
1. Connect to pi on CRYRING local network
ssh pi@192.168.207.127
Password available on request
2. To check state of baking code
screen -x
please do not press any keyboard button if unsure how the code works.
Pressing q instantly stops baking in case of emergencies
when done, to disconnect
CTRL+A
d
3. To check log of commands / previous status of baking code
cd Programs/Baking
cat Log.dat
4. To check temperature plot
cd Programs/Baking
gnuplot Plot_temps.gp
CTRL+C on terminal to quit
Please remember the Pi has limited computing power and the gnuplot window will consume some of that when open
One remote user is no problem, but multiple may give issues to the Pi
5. Check current pressure
Currently:
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?e=hEoZt7
(Pictures from webcam are in ~/Webcam on the pi
CM52 data also to be available online in real time soon) |
|
76
|
Thu Sep 23 13:31:12 2021 |
JM RSS ML | IE514 Gauge and Bakeout |
CARME setup checked by vaccum group this morning.
Both IE514 gauges de-gassed. To de-gas on the CM52 controller select the channel and hold the right arrow for two seconds. The reading will flash indicating the de-gassing procedure, this lasts for 2 minutes before completing. Flashing automatically stops once procedure finishes. Attached for convinience is the wire colour-pin assignment for the gauge head again.
Baking jacket fully assembled around CARME and baking code updated. Temperature and rate limits updated to match thermocouples around the chamber.
Max Temperature Max Rate
Chamber 155C 10C
Floating 175C 150C
Octagon (outside tent) 50C 10C
Automatic baking started at 13:50, pressure was 1.69e-7 mbar |
|
75
|
Wed Sep 22 14:44:51 2021 |
JM RSS ML | CARME leak test + bakeout tent |
On 21/9/21 CARME was connected to the leak tester. Leak testing showed no apparent leaks, extra time required to reach <1e-4 mbar during initial turbo pumpdown might be due to outgassing from additional installed components on the chamber. Turbo pump restarted
21/9/21 16:28 9.73e-7 mbar
21/9/21 16:48 8.92e-7 mbar
21/9/21 17:28 7.88e-7 mbar
elog 74 contains the pressure taken every hour after this point from the webcam.
NEG and NEG-ion cables for controller modules connected to feedthroughs on the CARME chamber. Feedthroughs have been electrically checked and are OK.
CARME baking jacket constructed around chamber. Thermocouples in place around the chamber. Vacuum group are coming tomorrow morning to inspect and de-gas installed gauge, after they are finished we mount final pieces of baking jacket and begin baking.
|
|
74
|
Wed Sep 22 08:30:08 2021 |
CB (remotely) | CARME pressure during first XHV test |
CARME pressure moved to
https://uoe-my.sharepoint.com/:x:/g/personal/cbruno2_ed_ac_uk/EVbsVOubNHRLj0PKiFIvd00BClMJ6WSJ1UDAyLA2F3R4jA?e=hEoZt7
Some photos overnight are over-exposed and pressure cannot be read.
Might be able to fix with image processing if needed, but likely not worth the time.
Baking started: 23/09/2021 14:00
Photo frequency changed to once per 10 minutes. |
|
73
|
Tue Sep 21 08:20:39 2021 |
ML | CARME installations |
Quote: |
Ion pump, RGA and additional gauge installed on the chamber.
Our gauge is the one mounted above the NEG feedthrough, Michael's gauge is mounted below the feedthrough. Ion pump too large to install on the top of the vessel, sufficient room on flange immediately next to this top flange.
Baking frame assembled around chamber, all required components fit within the baking volume. Insulation on baking tray and heaters set up and connected to pi controller.
Pumpdown begun at 16:30. Turbopump on at 16:40. Unable to reach the required pressure for the hot cathod gauge before leaving the lab ~18:10. Suggests possible leak, pressure and leak check tomorrow morning.
MACOR strain relief and plastic strain relief tested in edinburgh compared (see image). MACRO hole size is 2x bigger than the test piece, suggests mix up between radius and diameter for MACOR strain reliefs.
|
Vacuum group confirmed temperature limits for the RGA of 200°C without the electronics removed and 80°C with electronics installed. |
|
72
|
Mon Sep 20 11:45:45 2021 |
JM RSS ML | CARME installations |
Ion pump, RGA and additional gauge installed on the chamber.
Our gauge is the one mounted above the NEG feedthrough, Michael's gauge is mounted below the feedthrough. Ion pump too large to install on the top of the vessel, sufficient room on flange immediately next to this top flange.
Baking frame assembled around chamber, all required components fit within the baking volume. Insulation on baking tray and heaters set up and connected to pi controller.
Pumpdown begun at 16:30. Turbopump on at 16:40. Unable to reach the required pressure for the hot cathod gauge before leaving the lab ~18:10. Suggests possible leak, pressure and leak check tomorrow morning.
MACOR strain relief and plastic strain relief tested in edinburgh compared (see image). MACRO hole size is 2x bigger than the test piece, suggests mix up between radius and diameter for MACOR strain reliefs. |
|
71
|
Sun Sep 19 09:11:34 2021 |
CB (remotely) | Baking RasPi status |
Accessed Pi remotely
( 1. Without the Internet, the Pi has no way of knowing what time is it.
In case of power loss make sure to set correct date via
sudo date -s 'YYYY-MM-DD HH:MM:SS'
Or all log files will have wrong timestamp )
EDIT: Pi now connected to GSI time server.
2. Temperature logging OK
3. Commands/events logging enabled. Logging to file Log.dat, automatically moved to Logs/$(timestamp)_log.dat on
gracefuly quit.
Cannot test due to TC08 / RE72 not being connected. Please test when possible.
4. Plot_temps.gp gnuplot script currently optimised for 1080p monitor. Monitor in use is 1024x768.
Please optimise first line of script (size, position) to make it fit while we search for a better monitor.
5. Hostname changed to CARMEbakingpi |
|
70
|
Fri Sep 17 16:42:25 2021 |
JM RSS ML + OG | CARME Realignment + Thermocouple installation |
Laser surveying team worked till 11. Chamber was too high by 3mm close to the jet target and 6mm by the turbo pump (towards the dipole).
CARME lowered appropriate amount using bolts under CARME feet. Height checked using a laser on the centre point line of the middle flange turbo end and mid point on the front of the chamber (see images). During lowering front foot swapped and valve supports moved slightly to give more clearance for opening. ~13cm clearance for opening the chamber currently.
Pipe and single bellow connecting CARME to dipole replaced by two bellows and new pipe. Rest of open ports on chamber closed.
Thermocouple feedthrough flange installed under the motors flange in the turbo pump section. One internal thermocouple hooked onto kapton cable connected to MACOR and one thermocouple left floating in the chamber. Floating thermocouple corresponds to right output on feedthrough, thermocouple on kapton corresponds to left output on feedthrough. Image taken from flanges at the top of the chamber, quality is poor due to lack of maneouvering space.
Baking frame partially assembled around the chamber. Full assembly requires removing valve airlines to get under the valve pipe, will be done Monday morning. Took measurements for the RGA around the chamber, should be able to fit in multiple positions without conflict with baking jacket. |
|
69
|
Thu Sep 16 15:24:29 2021 |
JM RSS | CARME vented + thermocouple feedthrough |
Chamber has been vented back to atmosphere.
Motor was able to be pulled back to insert bloacking pin once at atmosphere.
Thermocouple feedthrough constructed with 4 thermocuples attached. Internal thermocouples each ~200mm in length with external thermocouples each ~1m in length. All were tested by plugging into TC08 and using baking code. Feedthrough will be installed under the motors flange tomorrow.
Survey team come tomorrow morning, afterwhich ion pump and RGA will also be installed. |
|
68
|
Sat Sep 11 16:01:44 2021 |
CB RSS | Preparations for shutdown |
Found CARME pressure at 2E-7 mbar
Raspberry pi took pictures until 4 am and
then lost contact with camera. Reasons
unclear. Plugging camera in / out solved
the issue. Will have to be investigated
if it happens again.
Left list of things to do with RSS. In
particular:
- Install xhv feedthrough and
thermocouples. Prepared on desk
- Install blocking pin in right hand motor
after pressure returns to atmosphere
- Cover bellows with aluminium to protect
from fibreglass when mounting tent
eventually
- Check all parts available for detector
supports with mounting sketch Peter sent (
RSS to recover sketch)
- Investigate baking program logging.
Unclear if logging is working - did we
disable it for some reason?
Valve on scroll closed. Scroll off. Turbo
off. Pressure sensors off. Raspberry pi
off.
Activities to resume on 16 September.
Attached are the pumping down plots
start point: 10.09.2021 19:00
end point: 11.09.2021 16:00
10.09.2021 19:00 1.09e-6
10.09.2021 20:00 7.57e-7
10.09.2021 21:00 6.17e-7
10.09.2021 22:00 5.36e-7
10.09.2021 23:00 4.82e-7
11.09.2021 00:00 4.42e-7
11.09.2021 01:00 4.10e-7
11.09.2021 02:00 3.84e-7
11.09.2021 03:00 3.62e-7
11.09.2021 04:00 3.43e-7
11.09.2021 15:00 2.29e-7
11.09.2021 16:00 2.31e-7 |
|
67
|
Fri Sep 10 18:09:14 2021 |
CB RSS ML | Leak test + baking system test |
Leak tested CARME via He leak tester + He
bottle. No leaks could be found. Some
indication of potential leak close to
blind flange on bottom valve - disappeared
after closer checks.
Outgassing from cables? If so second turbo
may be required for baking.
Removed leak tester. Scroll on, turbo on.
Pumping down overnight.
Tested baking system. Missing:
- Monitor (not shipped from DL??)
- Monitor power cable
- Power cable for USB hub (not shipped??)
Could find the first two, but monitor is
of poor quality.
USB to RS232 cable damaged in shipping
(poorly secured). Still works, may be
worth repairing but not trivial. Only
works plugged in to top USB 3 port on Pi.
System tested and still working aside from
these issues.
Connected webcam to pi facing IE514
display. Will take pictures every hour on
the hour as in DL.
Will stop pumping tomorrow pm due to
imminent shutdown.
Pressure at 5E-6 mbar decreasing. |
|
66
|
Fri Sep 10 10:50:55 2021 |
CB RSS ML | Pressure after overnight pump down |
Pressure found in 10-7 mbar range.
Decided to leak test anyway for safety.
Closed top valve. Vented turbo and backing
line. Installed additional valve and T
piece.
Pumping down again. Will begin leak
testing soon. |
|
65
|
Thu Sep 9 13:01:07 2021 |
CB RSS + OG | All feedthrough flanges mounted and motor rotated |
Moved cables that touched neg pumps. See
photos. Please bear in mind perspective.
Note avoiding touching chamber
bottom is not possible.
Re-installed motor to fix alignment
issues. All good. See attached.
Started pumping down. Reached 3E-1 mbar in
20 minutes (ok). Turbo on. After one hour
turbo stable at 0.85 A, pressure on Pirani
at 1.1E-2 (not reliable).
IE514 doesn't report pressure, i.e. > 1E-4
mbar.
Leak likely. Will try to pinpoint
tomorrow. We only changed 10 flangese. |
|
64
|
Wed Sep 8 17:03:07 2021 |
CB RSS + OG NB | Six feedthrough flanges mounted |
Continued to mount flanges as per previous ELOG entry.
Peter also remarked that the wires look a bit loose inside the holes, although this is how they were during
dirty tests in Edinburgh.
Attempted to pack the wires in two holes rather than 4 - too tight. Cannot close.
Packed the wires in 3 holes instead. Better, but still a bit loose. Wire harness feels quite robust and no
obvious strain is placed on the connectors.
#1 flange left with four strands for now.
Connectors were tested in this way in Edinburgh for motion tests. No broken cables after several thousand
repetitions.
Had troubles with second flange being mounted. Glove was caught under MACOR connector. Managed to free it
somehow with a cleaned screwdriver. Other half of the flange refused to go in due to bent pins, until NB was
able to repair it.
Mounted as #5.
Most other connector pins OK. Had to adjust a very small number. Cause of damage unclear - transportation? See
photos.
Took some pictures inside the chamber. Most connectors look OK, but
- One harness is touching the NEG wires (#2). This could be an issue. Will try to move it tomorrow
- Wires hanging from the top may be putting a lot of weight on the connectors. See pictures. Given number of
wires this may be acceptable.
- One harness is touching the NEG cage. Definitely an issue. Will fix tomorrow
Top/bottom flanges mounted with short harnesses. Side flanges mounted with long harnesses.
Two flanges left to mount.
Only #1 has been tightened. Other flanges left finger tight for Grossmontage to tighten properly tomorrow.
Outside motor to be fixed tomorrow. |
|
63
|
Wed Sep 8 10:52:54 2021 |
CB RSS + OG | First feedthrough flange mounted |
Assembled feedthrough flange with re-cut
SS support bracket. M3 vented on the
sides, M3x30 partly threaded on top. Top
screws don't touch the
Macor (by design? See first attachment).
Macor affixed but has
some play - good for thermal dilation.
Mounted long wire harness (side flange) on
top. Unable to get one of the two
connectors fully in. Unable to get it out
as well. See last attach, it's a small gap.
Tested few pins with multimeter - all ok.
Decided to mount.
No issues mounting. Slowly and carefully
threaded wires harness in CARME taking
care not to hit negs.
Slightly concerned about weight of Macor
connector on wires, but no obvious
alternative.
Please have a look at attachments and
comment. Will keep mounting in the
afternoon. |
|
62
|
Tue Sep 7 14:10:02 2021 |
CB RSS + OG | Strain relief SS support re-drilled |
Re-drilled SS supports in order to fit M3 screws.
Used hand tool cleaned with alchool for 6 of the 8.
Had to use machine for another two with thread damaged.
All 8 will be sent for cleaning. Should be able to start mounting at least the 6 ones tomorrow.
Reason for lack of fit still not clear. They were tested in Edinburgh. Possibly result of high-temperature baking? |
|
61
|
Tue Sep 7 12:40:43 2021 |
CB ML + RH | CARME survey preparation |
Rudi and survey group representative came
to CARME.
Elephant feet confirmed to be out of
position with respect to the models. Still
unclear if CARME is where it should be
with respect to the jet target and
electron target. I.e. do they have space
to mount? Do we have to move CARME again
(hopefully not!!)
Valve to downstream confirmed out of
alignment. Solution unclear but may be
needed to ensure beam goes through
detectors correctly.
Survey group will come to confirm CARME
position on Thursday 16 September. We
cannot mount baking tent until then, but
no major issue given delays due to strain
relieves and shutdown on 12-15 September. |